### v1

@@ Abstract
In this work, we investigate the spike-adding mechanism in a class of three-dimensional fast-slow systems with three distinct timescales, inspired by the FitzHugh-Nagumo (FHN) model driven by periodic input. First, we numerically generate a bifurcation diagram for the FHN model by varying the frequency and amplitude of the input, revealing that as the frequency decreases and the amplitude increases, the number of spikes within each burst grows. Next, we apply methods from geometric singular perturbation theory to compute critical and super-critical manifolds of the fast-slow system. We use them to characterize the emergence of new burst-spikes in the FHN model, when the periodic forcing resembles a low frequency-band brain rhythm. We then describe how the uncovered spike-adding mechanism defines the boundaries that separate regions with different spike counts in the parameter space.
keywords:
Spike-adding, parabolic bursting, circle/circle bursting, folded saddle canard, folded node canard, delayed-Hopf.
{MSCcodes}
34C23, 34C25, 34E13, 34D15, 92B05, 37N25.
@@ 1 Introduction
A common firing pattern in neurons and other excitable cells is bursting, which is characterized by an active phase (a sequence of spikes) followed by a silent phase. Early work by Rinzel and collaborators [31, 32, 33] pioneered the classification of bursting oscillations by fast/slow geometric methods, focusing on the bifurcation of the fast subsystem where the slow variables are treated as fixed parameters. Initially, they identified three classes of bursts: square-wave, parabolic, and elliptic. Later, in his classification scheme [23], Izhikevich renamed these bursting types based on the bifurcations of the corresponding equilibrium and limit cycles. For example, square-wave, parabolic and elliptic bursting were renamed fold/homoclinic, subHopf/fold cycle and circle/circle bursting, respectively.
@@
Two key questions regarding bursts, which were not addressed comprehensively by previous classifications, concern the transition from spiking to bursting and the process of adding spikes within a burst. While the former question has received considerable attention [17, 20, 22, 29, 34, 36, 38], the latter remained less explored. In particular, the spike-adding mechanism in square-wave bursting has been studied both numerically and analytically through fold/homoclinic bifurcations [2, 3, 4], codimension one- and two- homoclinic bifurcations [11, 28], and via canard explosions [10, 29, 30]. In contrast, there has been limited research on spike-adding mechanisms of parabolic and elliptic bursts. One example is Desroches et al.’s paper [16] that analyzed parabolic bursting in a system with 2-slow/2-fast dynamics. Their four-dimensional system had folded saddle canards that were shown to define the mechanism for burst spike-generation.
@@
In our paper we present novel results for the parabolic bursting by studying how it forms and how it increases its number of spikes in a three-dimensional system with three distinct timescales. As a toy-model for our analysis we consider the FitzHugh-Nagumo (FHN) system with the fast-timescale equation (“the voltage equation”) driven by a low-frequency periodic input. First, we investigate numerically the dynamics of the periodically forced FHN equations as the amplitude and frequency of the forcing term vary. We find a transition in the system’s dynamics from single-spike to bursting as the input frequency decreases and its amplitude increases. We then focus on the slowly oscillating force-drive, where the system exhibits bursting, and describe the underlying spike-adding mechanism. The forced FHN model can be converted to a three-dimensional system with three timescales (fast/slow/super-slow) and exhibits parabolic bursting. We show that the bursting starts and terminates at saddle-node bifurcations on invariant circles (or circle/circle bifurcations, as they are sometimes referred to; [23]).
@@
Other studies [14, 27] have discussed a similar problem in three-dimensional systems with three time scales. In [27], the spike-adding mechanism is attributed to the presence of a delayed Hopf bifurcation and a folded node singularity, while in [14] it is due to a Hopf bifurcation and a folded saddle singularity. The dynamics analyzed in our paper are closely related to the setups in [14, 27]. However, they distinguish themselves by requiring a combination of the aforementioned mechanisms, with spike-adding due to both folded node and folded saddle singularities.
@@
The main contributions of this paper are threefold. We provide a detailed theoretical and numerical exploration of how the intrinsic dynamics of a neuronal model could be finely tuned by controlling the frequency and/or the amplitude of the external input. Specifically, the dynamics of interest are bursting patterns characterized by a given number of spikes within each burst. We also link our theoretical results to an explicit biological question: how patterns of spikes or bursts of action potentials recorded from neurons in vivo correlate with underlying brain rhythms like delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-12 Hz), beta (13-35 Hz) or gamma (35-70 Hz) [8, 37] or otherwise, how electrical stimulation modulates the number of spikes, and/or organize them in bursts, in neurons [1, 42]. Lastly, we contribute new results to the theory of parabolic bursting by identifying distinct dynamical mechanisms for it to emerge and grow or lose spikes in three-dimensional systems with three timescales.
@@
The paper is organized as follows. In Sec. 2, we introduce the model: a two-dimensional FHN system with periodic input. We analyze its dynamics numerically from a bifurcation perspective, as the frequency and amplitude of the input change, and provide a biological motivation for our study. In Sec. 3, we reformulate the slowly-driven two-dimensional FHN model into a three-dimensional fast/slow/super-slow system with three distinct timescales, and we investigate this system using geometric singular perturbation theory. We state our main results in Sec. 4: we show theoretically how spikes are added to or removed from bursts by linking the system’s trajectories to folded singularity equilibria. In Sec. 5, we describe how the spike-adding mechanisms identified in Sec. 4 define the boundaries in parameter space that separate regions with different spike counts. We conclude with a short discussion of our results and future directions in Sec. 6. Details of analytical calculations presented in this paper are included in the Appendix.
@@ 2 FHN model with slowly varying force drive
(1a) $\dfrac{dx}{dt}$ $=x-\frac{1}{3}x^{3}-y-a+E\sin(\omega t)$
(1b) $\dfrac{dy}{dt}$ $=\varepsilon(x-by)$
@@
For our choice of parameters $a$, $b$ and if we replace the input with a constant, $I(t)=E_{0}$, the FHN model (1) exhibits two subcritical Hopf bifurcations. Furthermore, near each of these bifurcation points the system has a saddle-node of limit cycles (SNLC) bifurcation (see e.g., [13]). For $E_{0}$ values smaller than the first Hopf point at $E_{01}$, the equilibrium of the system is stable and $x$ (“the voltage”) does not spike. As $E_{0}$ increases past $E_{01}$, the equilibrium becomes unstable and the system’s trajectories are attracted to a stable limit cycle, and $x$ starts spiking. When $E_{0}$ passes through the second Hopf point, $E_{02}$, the equilibrium point becomes stable once again, and $x$ becomes saturated. As shown in [13], the first SNLC bifurcation occurs for input amplitudes slightly smaller than $E_{01}$, at $E_{SN1}$, where a stable and an unstable limit cycle emerge. Then, as $E_{0}$ approaches the subcritical Hopf point, the amplitude of the unstable limit cycle decays to zero. On the other hand, the stable limit cycle increases its amplitude as $E_{0}$ approaches the Hopf point $E_{01}$, and persists until $E_{0}$ approaches the second SNLC bifurcation point at $E_{SN2}$. The parameter region $E_{SN1}<E_{0}<E_{01}$ where small amplitude limit cycles exist is narrow (exponentially small). The small amplitude limit cycles sharply transition to large amplitude relaxation oscillations in this narrow parameter space, a phenomenon known as a canard explosion. Solutions that spend non-negligible time along repelling regions of the phase space before making fast transitions to stable regions are called canard solutions. Since the canard solutions of the FHN model with constant input exist only in a narrow parameter space, the solutions are not robust and challenging to compute numerically. A similar result, but for $E_{02}<E_{0}<E_{SN2}$ is obtained near the upper subcritical Hopf point $E_{02}$.
@@
We computed numerically a bifurcation diagram for the FHN model with periodic input, and projected it on the parameter plane $(\omega,E)$; Figures 1-2 and Appendix 7.1 for simulation details. This bifurcation diagram depicts the boundaries between regions where the system’s bursting dynamics exhibit a different ratio $m:n$, where $m$ is the number of large amplitude oscillations, or spikes in $x$, and $n$ is the number of input periods ($T=2\pi/\omega$), as $\omega$ and $E$ vary. For example, in Figure 1, region I, the input frequency $\omega$ is very small ($\omega\ll\varepsilon$, where $\varepsilon=0.08$) and the system admits bursting as a sequence of spikes occurring over each input period (Figure 1; right panel I). In Figure 1, region II, $\omega$ falls in the range $\mathcal{O}(\varepsilon)$ to $\mathcal{O}(\sqrt{\varepsilon})$ as $\sqrt{\varepsilon}\approx 0.3$, and the system exhibits tonic spiking, with one spike over each input period. Then, in region III, where the input frequency $\omega$ is between $\mathcal{O}(\sqrt{\varepsilon})$ and $\mathcal{O}(1)$, the system exhibits mixed mode oscillations: a mixture of $m<n$ spikes and $n-m$ small amplitude oscillations occurs over $n$ input periods (Figure 1, right panel III: the occurrence of one spike and one small amplitude oscillation over two input periods). Lastly, in region IV, where $\omega$ is large ($\omega=\mathcal{O}(1)$), no spikes occur and only small amplitude oscillations may exist. In summary, the FHN model with periodic drive exhibits busting dynamics, and the number of spikes in each burst increases as $E$ increases and $\omega$ decreases (see also Figure 2).
@@
Figure 1: Left panel: The bifurcation diagram of FHN system (1) is shown as the frequency $\omega$ and amplitude $E$ of the periodic input vary. Four regions (from left to right) indicate four different behaviors of the system in terms of its number of spikes ($m$) versus the number of input periods ($n$). (I) The system exhibits bursting, i.e., $m>n$. This region is plotted in more detail in Figure 2; (II) The system exhibits tonic spikes, i.e., $m=n$; (III) The system exhibits mixed mode oscillations, i.e., $m<n$; (IV) No spikes occur ($m=0$). Right panel: A trajectory of each region from the left panel is shown over two periods of the input ($n=2$) and $m=6,2,1$, and 00, respectively. The ticks at the top of the figure depict possible values $\omega=\varepsilon=0.08$, $\omega=\sqrt{\varepsilon}\approx 0.3$, $\omega=1$, given that $\varepsilon=0.08$ in system (1).
@@
In this paper, we extend the numerical results shown in Figures 1 and 2, to a theoretical study of the FHN model (1) with slowly varying force drive. That means, we will focus our analysis on region I from Figure 1 where the frequency of the input is very small, $\omega\ll\varepsilon$, say $\omega=\mathcal{O}(\varepsilon\sqrt{\varepsilon})$ or lower. In this region, for a fixed value of $\omega$, over each input period $n=1$, the system could generate either no spikes ($m=0$), or one spike ($m=1$) or a burst with multiple spikes ($m>1$); see Figure 2, regions I through IV.
@@
Figure 2: Left panel: An enhancement of the blue region from Figure 1, which marks the spike-adding bifurcation diagram of FHN system (1) as the frequency $\omega$ changes between 0 and $\varepsilon$ and amplitude $E$ varies between 0.3 and 0.7. For each input period, the system can exhibit no spikes ($m=0$, light-gray), one spike ($m=1$, gray), or a burst with multiple spikes ($m>1$, diverse shades of blue). The squares indicate parameter values for exemplary trajectories. Right panel: Solutions of (1), with $(\omega,E)$ indicated by I–IV on the left panel, plotted over one period of the input ($n=1$). Shown here is an increasing sequence of $E$ values with $\omega$ fixed that yields increasing spike counts. Note that $\varepsilon=0.08$ in system (1).
@@
The FHN model is a two-dimensional reduction of the Hodgkin-Huxley model for neuronal activity that reproduces well the neuron’s excitability (activation/deactivation dynamics), and the large excursions in the neuron’s trajectory called spikes [24]; Figure 2, right panels II-IV. The original equations depicting the dynamics of the neuron’s membrane potential, $dV/dt=V-V^{3}/3-W+I(t)$, and a recovery variable, $dW/dt=0.08(V+0.7-0.8W)$, are equivalent to system (1) by taking $x=V$ and $y=W-a$ for $a=0.875$. The independent variable $t$ in (1), while unitless, measures the physical time in milliseconds. Therefore, the angular frequency in input $I(t)=E\sin(\omega t)$ gets calculated from the true frequency $f$ (in Hz, or $\mbox{s}^{-1}$) as $\omega=2\pi f/1000$. Accordingly, diverse frequency bands for possible stimulation, or brain rhythms recorded simultaneously with neuron spike trains such as delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-12 Hz), beta (13-35 Hz), gamma (35-70 Hz) and high-gamma (70-200 Hz) would correspond to $\omega$ in the orders of magnitude shown in Figures 1-2. Specifically, frequencies $0.5$ Hz, $4$ Hz, $8$ Hz, $12$ Hz, $35$ Hz, $70$ Hz, $200$ Hz define $\omega\approx 0.003$, $0.025$, $0.05$, $0.075$ (Figure 2) and $0.22$, $0.44$ and $1.26$ (Figure 1) respectively. At $\varepsilon=0.08$, these values show that our analysis has biological relevance. Indeed, by taking $\omega$ in the parameter range $\mathcal{O}(\varepsilon^{2})-\mathcal{O}(\varepsilon\sqrt{\varepsilon})-\mathcal{O}(\varepsilon\sqrt[4]{\varepsilon})$ we provide a mechanistic explanation on how spikes are added or subtracted from neuronal busting patterns, when they are stimulated in the low-frequency but physiological meaningful bands delta and theta (Figure 2, in different shades of blue). We investigate this problem in the next sections.

### v2

@@ Abstract
In this work, we investigate the spike-adding mechanism in a class of three-dimensional fast-slow systems with three distinct timescales, inspired by the FitzHugh-Nagumo (FHN) model driven by periodic input. First, we numerically generate a spike-adding diagram for the FHN model by varying the frequency and amplitude of the input, revealing that as the frequency decreases and the amplitude increases, the number of spikes within each burst grows. We demonstrate that a similar spike-adding structure occurs in the more realistic, periodically forced Morris–Lecar neuronal model. Next, we apply methods from geometric singular perturbation theory to compute critical and super-critical manifolds of the fast-slow system. We use them to characterize the emergence of new burst-spikes in the FHN model, when the periodic forcing resembles a low frequency-band brain rhythm. We then describe how the uncovered spike-adding mechanism defines the boundaries that separate regions with different spike counts in the parameter space.
Keywords Spike-adding diagram, folded saddle canard, folded node canard.
MSCcodes 34C23, 34C25, 34E13, 34D15, 92B05, 37N25.
@@ 1 Introduction
A common firing pattern in neurons and other excitable cells is bursting, which is characterized by an active phase (a sequence of spikes) followed by a silent phase. Early work by Rinzel and collaborators [1, 2, 3] pioneered the classification of bursting oscillations by fast/slow geometric methods, focusing on the bifurcation of the fast subsystem where the slow variables are treated as fixed parameters. Initially, they identified three classes of bursts: square-wave, parabolic, and elliptic. Later, in his classification scheme [4], Izhikevich renamed these bursting types based on the bifurcations of the corresponding equilibrium and limit cycles. For example, square-wave, parabolic and elliptic bursting were renamed fold/homoclinic, circle/circle, and subHopf/fold cycle bursting, respectively.
@@
Two key questions regarding bursts, which were not addressed comprehensively by previous classifications, concern the transition from spiking to bursting and the process of adding spikes within a burst. While the former question has received considerable attention [5, 6, 7, 8, 9, 10, 11, 12, 13], the latter remained less explored. In particular, the spike-adding mechanism in square-wave bursting has been studied both numerically and analytically through fold/homoclinic bifurcations [14, 15, 16], codimension one- and two- homoclinic bifurcations [17, 18], and via canard explosions [19, 8, 20]. In contrast, there has been limited research on spike-adding mechanisms of parabolic and elliptic bursts. One example is Desroches et al.’s paper [21] that analyzed parabolic bursting in a system with 2-slow/2-fast dynamics. Their four-dimensional system had folded saddle canards that were shown to define the mechanism for burst spike-generation.
@@
In this paper, we study the spike-adding structure of a class of bursting generated by a three-dimensional system with three distinct timescales. While this bursting does not necessarily fall into any of the three main classification categories, it does share some features with the elliptic bursters, and others with the parabolic bursters (see Discussion). As a toy-model for our analysis we consider the FitzHugh-Nagumo (FHN) system with the fast-timescale equation (“the voltage equation”) driven by a low-frequency periodic input. First, we investigate numerically the dynamics of the periodically forced FHN equations as the amplitude and frequency of the forcing term vary. We find a transition in the system’s dynamics from single-spike to bursting as the input frequency decreases and its amplitude increases. Next, by converting the forced FHN model to a three-dimensional system with three timescales (fast/slow/super-slow), we focus on the slowly oscillating force-drive, where the system exhibits bursting, and describe the underlying spike-adding mechanism. To emphasize the biological relevance of our mathematical results, we then demonstrate that a similar spike-adding structure occurs in the more realistic Morris–Lecar (ML) neuronal model. We choose parameters for the unforced ML model that produce two subcritical Hopf bifurcations and nearby saddle-node of limit cycles [22], similar to the unforced FHN model.
@@
Other studies [23, 24] have discussed a similar problem in three-dimensional systems with three time scales. In [24], the spike-adding mechanism is attributed to the presence of a delayed Hopf bifurcation and a folded node singularity, while in [23] it is due to a Hopf bifurcation and a folded saddle singularity. The dynamics analyzed in our paper are closely related to the setups in [23, 24]. However, they distinguish themselves by requiring a combination of the aforementioned mechanisms, with spike-adding due to both folded node and folded saddle singularities. Furthermore, our study focuses on a parameter region where the interaction between the folded singularities and the existing delayed Hopf is not apparent. In contrast, the interplay between the dynamics at the folded node and the folded saddle influences the bursting structure.
@@
The main contributions of this paper are threefold. We provide a detailed theoretical and numerical exploration of how the intrinsic dynamics of a neuronal model could be finely tuned by controlling the frequency and/or the amplitude of the external input. Specifically, the dynamics of interest are bursting patterns characterized by a given number of spikes within each burst. We also link our theoretical results to an explicit biological question: how patterns of spikes or bursts of action potentials recorded from neurons in vivo correlate with underlying brain rhythms like delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-12 Hz), beta (13-35 Hz) or gamma (35-70 Hz) [25, 26] or otherwise, how electrical stimulation modulates the number of spikes, and/or organize them in bursts, in neurons [27, 28]. Lastly, we contribute new results to the general theory of bursting by identifying distinct dynamical mechanisms for burst patterns to emerge and gain or lose spikes in three-dimensional systems with three timescales.
@@
The paper is organized as follows. In Sec. 2, we introduce the model: a two-dimensional FHN system with periodic input. We analyze its dynamics numerically by constructing spike-adding diagrams that show how the number of spikes per input cycle varies with the frequency and amplitude of the input. We also provide a biological motivation for our study. In Sec. 3, we reformulate the slowly-driven two-dimensional FHN model into a three-dimensional fast/slow/super-slow system with three distinct timescales, and we investigate this system using geometric singular perturbation theory. We state our main results in Sec. 4: we show theoretically how spikes are added to or removed from bursts by linking the system’s trajectories to folded singularity equilibria. In Sec. 5, we describe how the spike-adding mechanisms identified in Sec. 4 define the boundaries in parameter space that separate regions with different spike counts. In Sec. 6, we demonstrate that a similar spike-adding structure occurs in the more realistic, periodically forced Morris–Lecar neuronal model. We conclude with a short discussion of our results and future directions in Sec. 7. Details of analytical calculations presented in this paper are included in the Appendix.
@@ 2 FHN model with slowly varying force drive
$\dfrac{dx}{dt}$ $=x-\frac{1}{3}x^{3}-y-a+E\sin(\omega t)$ (1a)
$\dfrac{dy}{dt}$ $=\varepsilon(x-by)$ (1b)
@@
For our choice of parameters $a$, $b$ and if we replace the input with a constant, $I(t)=E_{0}$, the FHN model (1) exhibits two subcritical Hopf bifurcations. Furthermore, near each of these bifurcation points the system has a saddle-node of limit cycles (SNLC) bifurcation (see e.g., [29]). For $E_{0}$ values smaller than the first Hopf point at $E_{01}$, the equilibrium of the system is stable and $x$ (“the voltage”) does not spike. As $E_{0}$ increases past $E_{01}$, the equilibrium becomes unstable and the system’s trajectories are attracted to a stable limit cycle, and $x$ starts spiking. When $E_{0}$ passes through the second Hopf point, $E_{02}$, the equilibrium point becomes stable once again, and $x$ becomes saturated. As shown in [29], the first SNLC bifurcation occurs for input amplitudes slightly smaller than $E_{01}$, at $E_{SN1}$, where a stable and an unstable limit cycle emerge. Then, as $E_{0}$ approaches the subcritical Hopf point, the amplitude of the unstable limit cycle decays to zero. On the other hand, the stable limit cycle increases its amplitude as $E_{0}$ approaches the Hopf point $E_{01}$, and persists until $E_{0}$ approaches the second SNLC bifurcation point at $E_{SN2}$. The parameter region $E_{SN1}<E_{0}<E_{01}$ is of order $\varepsilon$; within an exponentially small interval near $E_{SN1}$, small amplitude limit cycles exist. The small amplitude limit cycles sharply transition to large amplitude relaxation oscillations in this narrow parameter space, a phenomenon known as a canard explosion [19, 8, 20]. Solutions that spend non-negligible time along repelling regions of the phase space (i.e., near the repelling portion of slow manifolds) before making fast transitions to stable regions (the attracting portion of slow manifolds) are called canard solutions.
Since the canard solutions of the FHN model with constant input exist only in a narrow parameter space, the solutions are not robust and challenging to compute numerically. A similar result, but for $E_{02}<E_{0}<E_{SN2}$ is obtained near the upper subcritical Hopf point $E_{02}$.
@@
We computed numerically a spike-adding diagram for the FHN model with periodic input, and projected it on the parameter plane $(\omega,E)$; Figures 1-2 and Appendix 8.1 for simulation details. This spike-adding diagram depicts the boundaries between regions where the system’s bursting dynamics exhibit a different ratio $m:n$, where $m$ is the number of large amplitude oscillations, or spikes in $x$, and $n$ is the number of input periods ($T=2\pi/\omega$), as $\omega$ and $E$ vary. For example, in Figure 1, region A, the input frequency $\omega$ is very small ($\omega\ll\varepsilon$, where $\varepsilon=0.08$) and the system admits bursting as a sequence of spikes occurring over each input period (Figure 1; right panel A). In Figure 1, region B, $\omega$ falls in the range $\mathcal{O}(\varepsilon)$ to $\mathcal{O}(\sqrt{\varepsilon})$ as $\sqrt{\varepsilon}\approx 0.3$, and the system exhibits tonic spiking, with one spike over each input period. Then, in region C, where the input frequency $\omega$ is between $\mathcal{O}(\sqrt{\varepsilon})$ and $\mathcal{O}(1)$, the system exhibits mixed mode oscillations: a mixture of $m<n$ spikes and $n-m$ small amplitude oscillations occurs over $n$ input periods (Figure 1, right panel C: the occurrence of one spike and one small amplitude oscillation over two input periods). Lastly, in region D, where $\omega$ is large ($\omega=\mathcal{O}(1)$), no spikes occur and only small amplitude oscillations may exist. In summary, the FHN model with periodic drive exhibits busting dynamics, and the number of spikes in each burst increases as $E$ increases and $\omega$ decreases (see also Figure 2).
@@
Figure 1: Left panel: The spike-adding diagram of FHN system (1) is shown as the frequency $\omega$ and amplitude $E$ of the periodic input vary. Four regions (from left to right) indicate four different behaviors of the system in terms of its number of spikes ($m$) versus the number of input periods ($n$). A The system exhibits bursting, i.e., $m>n$. This region is plotted in more detail in Figure 2; B The system exhibits tonic spikes, i.e., $m=n$; C The system exhibits mixed mode oscillations, i.e., $m<n$; D No spikes occur ($m=0$). Right panel: A trajectory of each region from the left panel is shown over two periods of the input ($n=2$) and $m=6,2,1$, and 00, respectively. The ticks at the top of the figure depict possible values $\omega=\varepsilon=0.08$, $\omega=\sqrt{\varepsilon}\approx 0.3$, $\omega=1$, given that $\varepsilon=0.08$ in system (1).
@@
In this paper, we extend the numerical results shown in Figures 1 and 2, to a theoretical study of the FHN model (1) with slowly varying force drive. That means, we will focus our analysis on region A from Figure 1 where the frequency of the input is very small, $\omega\ll\varepsilon$, say $\omega=\mathcal{O}(\varepsilon\sqrt{\varepsilon})$ or lower. In this region, for a fixed value of $\omega$, over each input period $n=1$, the system could generate either no spikes ($m=0$), or one spike ($m=1$) or a burst with multiple spikes ($m>1$); see Figure 2, regions a through d.
@@
Figure 2: Left panel: An enhancement of the blue region from Figure 1, which marks the spike-adding diagram of FHN system (1) as the frequency $\omega$ changes between 0 and $\varepsilon$ and amplitude $E$ varies between 0.3 and 0.7. For each input period, the system can exhibit no spikes ($m=0$, light-gray), one spike ($m=1$, gray), or a burst with multiple spikes ($m>1$, diverse shades of blue). The squares indicate parameter values for exemplary trajectories. Right panel: Solutions of (1), with $(\omega,E)$ indicated by ad on the left panel, plotted over one period of the input ($n=1$). Shown here is an increasing sequence of $E$ values with $\omega$ fixed that yields increasing spike counts. Note that $\varepsilon=0.08$ in system (1).
@@
Brain rhythms across a wide range of frequency bands play critical roles in neural communication, cognitive processing, and brain health [39, 40]. Delta and theta oscillations are essential for sleep regulation and memory consolidation, particularly in the hippocampus, where theta rhythms coordinate encoding and retrieval processes [41]. Alpha and beta rhythms are associated with attentional control and sensorimotor integration, while gamma oscillations are implicated in perception, working memory, and conscious awareness [42, 39]. High-gamma activity, often phase-locked to slower rhythms like theta, reflects localized population spiking and is closely tied to high-level cognitive functions [43]. Disruptions across these bands are linked to various neuropsychiatric disorders, making them valuable biomarkers and potential targets for therapeutic interventions [44]. Together, these rhythms form a dynamic framework that supports efficient information processing in the brain.
The FHN model is a two-dimensional simplification of the Hodgkin-Huxley model for neuronal activity that reproduces well the neuron’s excitability (activation/ deactivation dynamics), and the large excursions in the neuron’s trajectory called spikes [45]; Figure 2, right panels bd. The original equations depicting the dynamics of the neuron’s membrane potential, $dV/dt=V-V^{3}/3-W+I(t)$, and a recovery variable, $dW/dt=0.08(V+0.7-0.8W)$, are equivalent to system (1) by taking $x=V$ and $y=W-a$ for $a=0.875$. The independent variable $t$ in (1), while unitless, measures the physical time in milliseconds. Therefore, the angular frequency in input $I(t)=E\sin(\omega t)$ gets calculated from the true frequency $f$ (in Hz, or $\mbox{s}^{-1}$) as $\omega=2\pi f/1000$. Accordingly, diverse frequency bands for possible stimulation, or brain rhythms recorded simultaneously with neuron spike trains such as delta (0.5-4 Hz), theta (4-8 Hz), alpha (8-12 Hz), beta (13-35 Hz), gamma (35-70 Hz) and high-gamma (70-200 Hz) would correspond to $\omega$ in the orders of magnitude shown in Figures 1-2. Specifically, frequencies $0.5$ Hz, $4$ Hz, $8$ Hz, $12$ Hz, $35$ Hz, $70$ Hz, $200$ Hz define $\omega\approx 0.003$, $0.025$, $0.05$, $0.075$ (Figure 2) and $0.22$, $0.44$ and $1.26$ (Figure 1) respectively. At $\varepsilon=0.08$, these values show that our analysis has biological relevance. Indeed, by taking $\omega$ in the parameter range $\mathcal{O}(\varepsilon^{2})-\mathcal{O}(\varepsilon\sqrt{\varepsilon})-\mathcal{O}(\varepsilon\sqrt[4]{\varepsilon})$ we provide a mechanistic explanation on how spikes are added or subtracted from neuronal busting patterns, when they are stimulated in the low-frequency but physiological meaningful bands delta and theta (Figure 2, in different shades of blue). We investigate this problem in the next sections. To further assess the biological relevance of our results, we also demonstrate that a similar spike-adding structure occurs in the more realistic, periodically forced Morris–Lecar neuronal model. See Section 6 for more details and comparison.